OpenStudy (anonymous):

Generate the first 5 terms of this sequence: (1 point) f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2. 1, 2, 3, 4, 5 1, 2, 3, 5, 7 1, 2, 3, 5, 8 1, 3, 5, 7, 9

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hi, Please help I'm stuck..

OpenStudy (misty1212):

you know \(f(1)\) adn \(f(2)\) right?

OpenStudy (misty1212):

what \(f(n) = f(n - 1) + f(n - 2), \) means is that the next number is what you get when you add up the last two

OpenStudy (misty1212):

so for example \(f(3)=f(2)+f(1)\)

OpenStudy (misty1212):

you got that?

OpenStudy (misty1212):

starts with \(1,2\) then the next one is \(1+2=3\)

OpenStudy (anonymous):

Wait what??

OpenStudy (anonymous):

@misty1212 I dont understand how to do more?

OpenStudy (misty1212):

add the previous two to get the next one

OpenStudy (anonymous):

To get any or just the next one?

OpenStudy (misty1212):

\[1,2,\overbrace{3}^{1+2}\]

OpenStudy (anonymous):

so then 2+3 ?

OpenStudy (misty1212):

so we are at \(1,2,3\) next will be yes that

OpenStudy (anonymous):

1,2,3,5,8?

OpenStudy (anonymous):

@misty1212 right?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

sure

OpenStudy (anonymous):

For the function f(x) = x + 7, what is the ordered pair for the point on the graph when x = 2b? (1 point) (2b, 2b + 7) (2b, x + 7) (x, x + 7) (x, 2b + 7)

OpenStudy (misty1212):

if \(x=2b\) then in the first slot for \((x,y)\) you need to see \(2b\)

OpenStudy (anonymous):

so 2b + 2b?

OpenStudy (misty1212):

?

OpenStudy (anonymous):

2b + y?

OpenStudy (misty1212):

\[f(x)=x+7\\ f(2b)=2b+7\]

OpenStudy (anonymous):

What?..

OpenStudy (misty1212):

lol i guess you are confused, lets go slow

OpenStudy (misty1212):

you have a function \(f(x)=x+7\) which means, take your number, whatever \(x\) is, and add \(7\) to it

OpenStudy (anonymous):

2b + 7?

OpenStudy (misty1212):

so for example \[f(3)=3+7=10\\ f(0)=0+7=7\\ f(-7)=-7+7=0\\ f(\pi)=\pi+7\\ f(\heartsuit)=\heartsuit+7\] and so on

OpenStudy (misty1212):

yes so \[f(2b)=2b+7\] is the second coordinate

OpenStudy (anonymous):

what would be the first?

OpenStudy (misty1212):

yes, a very very rare A

OpenStudy (anonymous):

A?

OpenStudy (misty1212):

yup

OpenStudy (anonymous):

OKay, i just have one more can you help me on this last one?

OpenStudy (misty1212):

lol sure